### v2

@@ 2.5.1 Evaluation Metrics
To evaluate the quality and performance of the large family of classification models, we computed several metrics including accuracy (proportion of correct predictions), F1-score (harmonic mean of precision and recall), and Matthews correlation coefficient (MCC), which is defined as:
MCC $=\frac{\text{TP}\times\text{TN}~{}-~{}\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}\,,$ (1)
where TP is the number of true positives, TN the number of true negatives, FP the number of false positives, and FN the number of false negatives. The MCC ranges from $-1$ to $1$ and on imbalanced datasets, it can be more informative than either accuracy or the F1-score [5]. Our imbalanced dataset includes 88 positive examples and 998 negative examples.
2.5.2 Variable Selection
@@
We performed variable selection on both the imputed and non-imputed datasets for the top-performing classification models used in this study. For each model described in Section 3.1, we performed a grid search with cross-validation on the hyperparameter space to determine the best scoring model. We then performed forward feature selection by beginning with zero features in our models and iterating through each available feature, computing each model’s score with that feature included in it. The feature resulting in the highest model score becomes a feature in the model, and the search for a second feature proceeds. The process continues until we have the desired number of features in the model or an additional feature does not significantly improve the model score. The results from variable selection are presented in Section 3.1.
@@ 3 Results
Table 2 presents the MCCs for the classification models explored in this study on the imputed and non-imputed datasets. Columns containing more than 400 NA’s were dropped before imputation and data cleaning, so that both the imputed and non-imputed datasets contained the same set of features. The top performing classifiers (highlighted in gray) were the multi-layer perceptron (MLPC), histogram-based gradient boosting (HGBC), and support vector machine (SVC), as each obtained an MCC score above or very close to 0.5 in the cross-validation and test for both imputed and non-imputed datasets with high significance. We note that AdaBoosting achieved similar performance to the declared top performing models, however we will not discuss it further as we will focus on HGBC, which uses the same underlying principle of an ensemble of weak learners but it is a more modern algorithm allowing additional hyperparameters to be set (e.g., regularization).
Models whose test MCC is very close to zero are no better than a random guess. Gaussian process, Naïve Bayes, QDA, and Decision tree classifiers all performed particularly poorly on our datasets. Although Gaussian process is analogous to SVM, its implementation used in this analysis is non-sparse, and hence the model uses all features for each example when computing the high-dimensional posterior. As such, a possible explanation for this model’s poor performance is that the posterior computed from the training examples is skewed. If the posterior is skewed, the Laplace approximation (which is centered on the mode) will not be accurate [15].
@@
Table 2: Best MCCs for all the classification models considered in this study. Top performing models (highlighted in gray) were multi-layer perceptron (MLPC), histogram-based gradient boosting (HGBC) and support vector machine (SVC), on datasets with both multiple imputation and without imputation. The results presented here are from models using all available features which have less than 400 NA’s.
@@
[…]
@@ 3.1 Top Performing Models
From among the large family of classification models considered (Section 2.5), three models performed consistently well with both imputed and and non-imputed data. They were the multi-layer perceptron classifier (MLPC), the histogram-based gradient boosting classifier (HGBC), and the support vector machine classifier (SVC). The results from these models without variable selection are highlighted in Table 2 in light gray, and their confusion matrices are presented in Figure 3. We explored forward feature selection on these three models to understand which variables are important in determining whether a potato clone is kept in the trials. The resulting MCC scores for the top three performing models using forward feature selection on the imputed and non-imputed datasets are presented in Table 3.
@@
Figure 3: Confusion matrices for the top performing models applied to both imputed and non-imputed datasets: multi-layer perceptron classifier (MLPC, called a Neural Net here) in first row, histogram-based gradient boosting classifier (HGBC) in middle row, and support vector machine (SVM) classifier in the last row. Variable selection was not performed here, however variables containing at least 400 NA’s were dropped before training. Numerical values within the confusion matrices represent prediction proportions.
Table 3: MCC for the top classification models after forward feature selection.
@@
Model $\Delta$CV MCC Test MCC CI Test MCC $\Delta$CV MCC Test MCC CI Test MCC
MLPC 0.476 0.608 (0.581, 0.633) 0.595 0.623 (0.584, 0.659)
HGBC 0.531 0.574 (0.546, 0.602) 0.511 0.645 (0.608, 0.680)
SVM 0.539 0.502 (0.471, 0.533) 0.731 0.623 (0.584, 0.659)
[…]
It is interesting to note that variable selection significantly improved the test MCC of MLPC on both imputed and non-imputed datasets, as well as the test MCC of HGBC on the non-imputed dataset. However, this form of variable selection seems to hinder the performance of the SVM classifier on the imputed and non-imputed datasets. One reason for this drop in performance could be that features were selected after the hyperparameter search. Both techniques use the SVM classifier, which in turn suppresses certain features through an $\ell_{2}$-norm penalty. Thus, the second application of SVM for feature selection would drop the variable with high penalty regardless of their overall contribution to the decision function.
Table 4 lists the variables that were selected for each of the three top-performing models on each imputed and non-imputed dataset. We highlight in bold the variables selected in all three models when applied to a particular dataset. On the imputed dataset the variables “Trial Region_COR”, “Trial Region_HER_Early”, “Yr in trial”, “%RB”, “Vascular”, “GDD 61–90” and “GDD 91–end” were selected for all three models. On the other hand, the common variables selected for the non-imputed dataset were “Trial Region_HER_Late”, “Yr in trial”, “Hollow Heart”, “Brown Center”, ‘Vascular”, “Shape Uniformity”, and “GDD 1–60” instead.
Two variables selected in the models using the imputed dataset, “Trial Region_COR” and “Trial Region_HER”, gained their significance due to only being present for the years 2013 and 2014. Respectively, they accounted for 4.9% and 5.2% of the total kept varieties for all regions and years present in the imputed dataset. This imbalance may be the sole reason these variables were selected as informative.
The variables selected by all three models and for both datasets were “Yr in Trial” and “Vascular”. The importance of “Yr in Trial” may be irrelevant as it indicates how long a particular potato variety survived selection during the three-year trial. “Vascular” refers to vascular discoloration, a darkening of the ring of tissue that surrounds the tuber just below its skin. The ring contains the tuber’s vascular system, and its discoloration can be indicative of bacterial or fungal infections or abiotic stress response. The vascular discoloration in tubers can lead to low processing quality and is an important trait for selection.
Table 4: Variables selected via forward feature selection (second column) using the corresponding top performing algorithm on each dataset (first column). Bold variable names highlight the variables that were selected in common between the algorithms when applied to the same dataset (imputed versus not imputed).
Model Variables Selected
MLPC (imputed dataset) “Trial Region_COR”, “Trial Region_HER_Early”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB”, “%RB 1s”, “%1s”, “4-6oz”, “Y2s + Over 20oz”, “Culls”, “Specific Gravity”, “Blackspot”, “IBS”, “Vascular”, “Tuber Shape”, “Eye Depth”, “Ctrl Ave”, “CA”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
HGBC (imputed dataset) “Trial Region_COR”, “Trial Region_HER”, “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_KF”, “Trial Region_ONT”, “Yr in trial”, “%RB”, “Over 20oz”, “4-6oz”, “Under 4oz”, “Culls”, “Tuber/Plant”, “Sugar Ends”, “Hollow Heart”, “Brown Center”, “IBS”, “Vascular”, “Russeting”, “Ctrl Ave”, “GDD 61–90”, “GDD 91–end”
SVM (imputed dataset) “Trial Region_COR”, “Trial Region_HER”, “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_KF”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB’’, “Y1s”, “%RB 1s”, “%1s”, “Y2s + Over 20oz”, “Specific Gravity”, “Fry Color Stem”, “Fry Color Bud”, “Sugar Ends”, “Blackspot”, “Vascular”, “Scab”, “GDD 61–90”, “GDD 91–end”
MLPC (non-imputed dataset) ‘Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_ONT”, “Yr in trial”, “%RB 1s”, “10-20oz”, “6-10oz”, “4-6oz”, “Length/Width”, “Specific Gravity”, “Fry Color Stem”, “Hollow Heart”, “Brown Center”, “Vascular”, “Tuber Shape”, “Shape Uniformity”, “Growth Cracks”, “Ctrl Ave”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
HGBC (non-imputed dataset) “Trial Region_HER_Early”, “Trial Region_HER_Late”, “Trial Region_ONT”, “Yr in trial”, “Total Yield”, “%RB”, “Y1s”, “%RB 1s”, “Over 20oz”, “10-20oz”, “Culls”, “Tuber/Plant”, “Hollow Heart”, “Brown Center”, “Blackspot”, “Vascular”, “Shape Uniformity”, “Eye Depth”, “Scab”, “Shatter Bruise”, “GDD 1–60”
SVM (non-imputed dataset) “Trial Region_HER_Late”, “Yr in trial”, “Total Yield”, “%RB”, “Y1s”, “%RB 1s”, “%1s”, “Tuber/Plant”, “Length/Width”, “Specific Gravity”, “Hollow Heart”, “Brown Center”, “Vascular”, “Skin Color”, “Tuber Shape”, “Shape Uniformity”, “Scab”, “Ctrl Ave”, “GDD 1–60”, “GDD 61–90”, “GDD 91–end”
[…]

### v3

@@ 2.5.1 Evaluation Metrics
To evaluate the quality and performance of the large family of classification models, we computed several metrics including accuracy, F1-score, and Matthews correlation coefficient (MCC). Each of these metrics provides unique insights into model performance, particularly in the context of imbalanced datasets.
Accuracy is defined as the proportion of correct predictions made by the model across all instances. It is calculated as:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (1)
where TP is the number of true positives, TN is the number of true negatives, FP is the number of false positives, and FN is the number of false negatives. While accuracy is intuitive and widely used, it can be misleading for imbalanced datasets where one class significantly outnumbers the other (Powers, 2011).
The F1-score is the harmonic mean of precision and recall, providing a balanced measure of a model’s performance. It is particularly useful when dealing with imbalanced datasets (Chicco and Jurman, 2020). The F1-score is calculated as:
[…]
where Precision = TP / (TP + FP) and Recall = TP / (TP + FN). The F1-score ranges from 0 to 1, with higher values indicating better performance.
The Matthews correlation coefficient (MCC) is a measure of the quality of binary classifications, taking into account true and false positives and negatives. It is defined as:
$\text{MCC}=\frac{\text{TP}\times\text{TN}-\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}$ (3)
The MCC ranges from $-1$ to $1$, where $1$ represents a perfect prediction, 00 indicates no better than random prediction, and $-1$ represents total disagreement between prediction and observation. On imbalanced datasets, MCC can be more informative than either accuracy or the F1-score as it provides a more balanced measure of the proportion of all classes predicted correctly (Chicco, 2017; Boughorbel et al., 2017).
Our dataset is significantly imbalanced, including 88 positive examples and 998 negative examples. In such cases, the use of MCC is particularly relevant as it takes into account the ratio of the confusion matrix size (Chicco and Jurman, 2020). This characteristic makes MCC especially suitable for evaluating binary classification performance on imbalanced datasets, providing a more reliable performance metric compared to accuracy or F1-score alone.
2.6 Variable Selection
@@
We performed variable selection on both the imputed and non-imputed datasets for the top-performing classification models used in this study. For each model described in Section 3.1, we performed a grid search with cross-validation on the hyperparameter space to determine the best scoring model. We then performed forward feature selection by beginning with zero features in our models and iterating through each available feature, computing each model’s score with that feature included in it. The feature resulting in the highest model score becomes a feature in the model, and the search for a second feature proceeds. The process continues until we have the desired number of features in the model or an additional feature does not significantly improve the model score. We validate the feature selection process using 5-fold cross-validation. For our study, we explore different fractions of features to select: 0.1, 0.3, 0.5, 0.7, and 0.9, and compare the respective results—we present them in Section 3.1.
2.7 Simulation Study
Simulation studies are essential for evaluating the performance of statistical and machine learning models under various conditions. The ADEMP (Aims, Data-generating mechanisms, Estimands, Methods, Perfomance measures) framework, as outlined by Morris et al. (2019), provides a structured approach to designing and reporting simulation studies. The present study leveraged the ADEMP framework to evaluate the generalizability of the top performing classification models (see 3.1) on simulated datasets that are independent from the original potato trials dataset.
2.7.1 Aims
The primary aim of this simulation study was to assess the performance of a subset of classification models used in the overall study, whose goal was to predict whether a potato variety will graduate to the next set of trials (according to binary response variable: true_keeps used in supervised training). We focused on the top three performing models, and added logistic regression (linear model) for baseline comparison. Focusing on these models helped us keep computation manageable and enabled us to use a comprehensive cross validated grid-search approach for each trial.
Specifically, we aimed to:
  1. 1.
    Evaluate the predictive accuracy of non-linear models (Neural Network, Histogram-based Gradient Boosting Classifier (HGBC), and Support Vector Machine (SVM)) compared to a linear model (Logistic Regression, LR).
  2. 2.
    Determine the best-performing model in terms of True Positives (TP), False Positives (FP), False Negatives (FN), True Negatives (TN), Area Under the Receiver Operating Characteristic Curve (ROC), and Matthew’s Correlation Coefficient (MCC).
  3. 3.
    Measure results variability between each simulated dataset by computing the standard error (SE) of their MCC scores.
  4. 4.
    Investigate how variations in data-generating mechanisms affect model performance.
  5. 5.
    Identify and explain both expected and unexpected effects of different data distributions on model performance.
2.7.2 Data-Generating Mechanisms
In simulation studies, the choice of data-generating mechanisms is crucial for evaluating the performance of statistical methods under various conditions. For example, Dasgupta et al. (2013) selected specific data-generating mechanisms in their study to comprehensively assess the robustness of different clustering methods. They used normal, binomial, Poisson, and Gamma distributions to reflect the diverse characteristics of their protein data and to test how clustering methods handle different distributional properties and data structures.
Similarly, in our study, we varied the data-generating distributions to simulate different scenarios. The original potato variety trials dataset (empirical dataset) contains measurements on potato clones, including weight (or mass) variables, counts, and ratings. These data types were taken into account in our simulation of different scenarios:
  1. 1.
    Scenario 1: $\mathrm{Normal}(\mu_{\text{data}},\sigma_{\text{data}})$ for weight/mass type quantities, $\mathrm{Poisson}(\mu_{\text{data}})$ for count type quantities, $\mathrm{Uniform}(a=\min_{\text{data}},b=\max_{\text{data}})$ for rating type quantities.
  2. 2.
    Scenario 2: $\mathrm{Normal}(\mu_{\text{data}},\sigma_{\text{data}})$ for weight/mass type quantities, $\mathrm{Poisson}(\mu_{\text{data}})$ for count type quantities, $\mathrm{Beta}(a=2,b=5)$ for rating type quantities.
  3. 3.
    Scenario 3: $\mathrm{Gamma}(k=2,\theta=2)$ for weight/mass type quantities, $\mathrm{Poisson}(\mu_{\text{data}})$ for count type quantities, $\mathrm{Uniform}(a=\min_{\text{data}},b=\max_{\text{data}})$ for rating type quantities.
  4. 4.
    Scenario 4: $\mathrm{Gamma}(k=2,\theta=2)$ for weight/mass type quantities, $\mathrm{Poisson}(\mu_{\text{data}})$ for count type quantities, $\mathrm{Beta}(a=2,b=5)$ for rating type quantities.
We applied normal and gamma distributions for weight/mass variables, Poisson distributions for count variables, and both uniform and beta distributions for rating variables across different scenarios. This approach ensured a thorough examination of the classification methods’ performance in handling different shapes and types of data, thereby providing a robust evaluation framework. The selection of these mechanisms allowed the capturing of inherent distributional characteristics present in the original dataset, thus enhancing the reliability and generalizability of the study’s findings.
For each scenario, 500 datasets were simulated to account for variability. Although a larger number of datasets would be best to fully confirm robustness, applying cross-validated parameter optimization gridsearch to each classification model incurs significant computational cost. Available computational resources limited our odyssey to 2001 datasets (4 $\times$ 500 simulated + original), each of 885 observations and 41 features. Due to the sample size, and in combination with the other methods employed, we believe this demonstrates a reasonable level of robustness. We leave the task of more intensive computational verification for future study and present our current insights.
2.7.3 Estimands
The primary estimands in the simulation study were:
[…]
  • The classification decision for each potato variety (i.e., predicted true_keeps values).
2.7.4 Methods in Simulation Study
Four classification models were used to predict the binary response variable true_keeps:
  1. 1.
    Multi-layer Perceptron Classifier (Neural Net)
  2. 2.
    Histogram-based Gradient Boosting Classifier (HGBC)
  3. 3.
    Support Vector Machine Classifier with RBF Kernel (SVM)
  4. 4.
    Logistic Regression (LR)
Each model underwent hyperparameter tuning via cross-validated grid search. Data imputation was performed, where the missingness from the original dataset was mimicked in the simulated datasets prior to imputation. All variables with no more than 400 missing values were considered under each model (i.e. there was no variable selection in the simulation study).
2.7.5 Performance Measures
The performance of the models was evaluated using the following metrics:
  • True Positives (TP)
  • False Positives (FP)
  • False Negatives (FN)
  • True Negatives (TN)
  • Area Under the Receiver Operating Characteristic Curve (ROC)
  • Matthew’s Correlation Coefficient (MCC)
  • Standard error (SE) between MCC scores
For each scenario, the metrics were averaged over the 500 simulated datasets.
@@ 3 Results
Table 2 presents the MCCs for the classification models explored in this study on the imputed and non-imputed datasets. Columns containing more than 400 NA’s were dropped before imputation and data cleaning, so that both the imputed and non-imputed datasets contained the same set of features. The top performing classifiers (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC), and support vector machine (SVC), as each obtained an MCC score above or very close to 0.5 in the cross-validation and test for both imputed and non-imputed datasets with high significance. We note that AdaBoosting achieved similar performance to the declared top performing models, however we will not discuss it further as we will focus on HGBC, which uses the same underlying principle of an ensemble of weak learners but it is a more modern algorithm allowing additional hyperparameters to be set (e.g., regularization).
Models whose test MCC is very close to zero are no better than a random guess. Gaussian process, Naïve Bayes, QDA, and Decision tree classifiers all performed particularly poorly on our datasets. Although Gaussian process is analogous to SVM, its implementation used in this analysis is non-sparse, and hence the model uses all features for each example when computing the high-dimensional posterior. As such, a possible explanation for this model’s poor performance is that the posterior computed from the training examples is skewed. If the posterior is skewed, the Laplace approximation (which is centered on the mode) will not be accurate (Rasmussen et al., 2006).
@@
Table 2: Best test Mathews correlation coefficients (MCC) and their respective 95% confidence intervals (CI) for all the classification models considered in this study. Top performing models (highlighted in gray) were the feedforward neural network classifier (Neural Net), histogram-based gradient boosting (HGBC) and support vector machine (SVC), on datasets with both multiple imputation and without imputation. The results presented here are from models using all available features which have less than 400 NA’s. QDA stands for Quadratic Discriminant Analysis.
@@
Based on the results presented in Table 2, if we had to pick three models, we should choose the Neural Net, Histogram-based Gradient Boosting (HGBC) and Support Vector Machine (SVM) models. These three models are highlighted in gray in the table as the top-performing models. Here’s the rationale for selecting these models:
Neural Net:
    Demonstrates strong performance on both datasets
    Has the second-highest MCC score (0.623) for the non-imputed dataset
    Shows improvement when moving from imputed (0.530) to non-imputed (0.623) data
HGBC:
    Performs well on the imputed dataset with an MCC of 0.563
    While its performance on the non-imputed dataset is lower (0.485), it’s still among the top performers overall
SVM:
    Performs exceptionally well on both imputed and non-imputed datasets
    Has the highest MCC score (0.692) for the non-imputed dataset
    Shows consistent performance across both datasets (0.563 for imputed, 0.692 for non-imputed)
These three models consistently outperform the others across both imputed and non-imputed datasets. They also offer a good variety of approaches (kernel-based, neural network, and ensemble method), which could be valuable for comparing and combining results in an ensemble approach.
It’s worth noting that AdaBoosting also performs well, especially on the non-imputed dataset (0.541). However, HGBC, which is a more advanced boosting technique, generally outperforms it, especially on the imputed dataset.
@@ 3.1 Top Performing Models
From among the large family of classification models considered (Section 2.5), three models performed consistently well according to their MCC scores, with both imputed and and non-imputed data. They were the feedforward neural network classifier (Neural Net), the histogram-based gradient boosting classifier (HGBC), and the support vector machine classifier (SVC). The results from these models without variable selection are highlighted in Table 2 in light gray, and their confusion matrices are presented in Figure 3.
The Neural Net model shows strong performance on both imputed and non-imputed datasets:
    High true negative rate (0.94) and true positive rate (0.60).
    Similar performance, with a slightly higher true positive rate (0.71) but lower true negative rate (0.96).
This suggests that the Neural Net is effective at both identifying varieties to keep and those to drop, with a slight improvement in identifying varieties to keep when using non-imputed data. The HGBC model demonstrates consistent performance across both datasets:
    High true negative rate (0.95) but lower true positive rate (0.67).
    Nearly identical performance to imputed data.
HGBC shows strong ability to identify varieties to drop but is less effective at identifying varieties to keep. The consistency between imputed and non-imputed datasets suggests that HGBC is robust to the imputation process. The SVM model shows interesting differences between imputed and non-imputed datasets:
    High true negative rate (0.96) and moderate true positive rate (0.61).
    Perfect true positive rate (1.00) but slightly lower true negative rate (0.96).
SVM demonstrates excellent performance, particularly on the non-imputed dataset where it perfectly identifies varieties to keep. However, this comes at a slight cost to its ability to identify varieties to drop.
Comparing the three models we can see that all models show strong performance in identifying varieties to drop (high true negative rates). SVM on non-imputed data shows the best performance for identifying varieties to keep, but may be overfitting. Neural Net and HGBC show more balanced performance across imputed and non-imputed datasets. Imputation appears to have a minimal impact on HGBC, moderate impact on Neural Net, and significant impact on SVM performance.
These results suggest that while all three models are effective for potato variety selection, the choice of model and whether to use imputation may depend on the specific goals of the breeding program (e.g., prioritizing identification of varieties to keep vs. varieties to drop).
We also explored forward feature selection on these three models to understand which variables are important in determining whether a potato clone is kept in the trials. The resulting accuracy, F1 and MCC scores for the top three performing models using forward feature selection with 10, 30, 50, 70 and 90 percent of features included on the imputed and non-imputed datasets are presented in Table 3.
Table 3: Performance metrics: Accuracy, F1-score, and Matthews Correlation Coefficient (MCC) for the top classification models after forward feature selection. All reported scores were statistically significant and within 95% confidence intervals.
% of Model Imputed Not Imputed
features Accuracy F1-score MCC Accuracy F1-score MCC
0.1 Neural Net 0.920 0.680 0.429 0.950 0.770 0.616
HGBC 0.850 0.640 0.300 0.930 0.780 0.564
SVM 0.930 0.740 0.513 0.920 0.480 0.000
0.3 Neural Net 0.890 0.710 0.423 0.930 0.720 0.440
HGBC 0.870 0.690 0.400 0.930 0.800 0.604
SVM 0.930 0.740 0.513 0.920 0.480 0.000
0.5 Neural Net 0.900 0.740 0.491 0.940 0.820 0.636
HGBC 0.890 0.720 0.443 0.960 0.860 0.716
SVM 0.900 0.730 0.472 0.920 0.480 0.000
0.7 Neural Net 0.900 0.730 0.472 0.960 0.840 0.697
HGBC 0.890 0.720 0.433 0.910 0.690 0.373
SVM 0.900 0.740 0.491 0.930 0.650 0.432
0.9 Neural Net 0.920 0.780 0.559 0.940 0.780 0.569
HGBC 0.910 0.740 0.488 0.930 0.720 0.440
SVM 0.900 0.760 0.536 0.940 0.720 0.531
@@
Figure 3: Confusion matrices for the top performing models applied to both imputed and non-imputed datasets: feedforward neural network classifier (Neural Net) in first row, histogram-based gradient boosting classifier (HGBC) in middle row, and support vector machine (SVM) classifier in the last row. Variable selection was not performed here, however variables containing at least 400 NA’s were dropped before training. Numerical values within the confusion matrices represent prediction proportions.
Overall, model performance generally improves as more features are included, with the best results often seen at 90% of features, or with no feature selection. The impact of imputation varies across models and metrics, suggesting that the effectiveness of imputation is model and data-dependent. Notably, no single model consistently outperforms the others across all scenarios, indicating that model selection should be context-specific.
The Neural Network shows consistent improvement with more features in both imputed and non-imputed datasets. It achieves the highest accuracy (0.96) and F1-score (0.84) on the non-imputed dataset at 70% feature selection, demonstrating robust performance across different feature selection stages.
The HGBC model exhibits variable performance across feature selection stages. It peaks in performance on the non-imputed dataset at 50% feature selection, with the highest accuracy (0.96) and F1-score (0.86) among all models, but shows more sensitivity to feature selection compared to other models.
SVM displays consistent performance in the imputed dataset across feature selection stages but notably underperforms in the non-imputed dataset for lower feature percentages, with MCC of 0 at 10%, 30%, and 50% feature selections. However, it improves significantly in the non-imputed dataset with higher feature percentages.
The impact of imputation varies among the models. The Neural Network generally benefits from imputation at lower feature percentages but shows comparable or better performance without imputation at higher percentages. HGBC shows mixed results, with imputation being beneficial in some cases and detrimental in others. SVM significantly benefits from imputation, especially at lower feature percentages where it struggles with non-imputed data.
As forward feature selection becomes more effective with almost no selection, it does not seem beneficial for these data wherein most features appear to contribute valuable information. That said, good performance can also be achieved at lower percentages of feature inclusion (e.g., 50% and 70%), indicating some potential for effective dimensionality reduction.
Not surprisingly, yield components emerged as the most important factors, with variables like ”Total Yield,” ”%RB,” ”Y1s,” ”%1s,” and ”Over 20oz” consistently selected across all models and datasets. Quality traits also played a substantial part in the classification process with features such as ”Specific Gravity,” ”Fry Color,” ”Hollow Heart,” and ”Brown Center” frequently chosen by the models. Environmental factors, particularly Growing Degree Days (GDD) variables, were also often included in the selected features.
Trial information proved to be another critical aspect, with ”Yr in trial” and ”Trial Region” consistently appearing in the selected variables. This indicates that the stage of the trial and the location where it’s conducted significantly influence the classification results. Interestingly, variables related to defects, such as ”Blackspot,” ”IBS,” and ”Vascular,” were more frequently selected in HGBC models, suggesting this particular algorithm may be more attuned to these quality issues.
The forward variable selection process showed notable consistency in the variables selected across imputed and non-imputed datasets. Many of the same variables were chosen regardless of the dataset type, indicating robustness in variable importance. However, there were some model-specific preferences, with different models prioritizing slightly different sets of variables. This variation could be attributed to the unique algorithms of each model and how they handle feature interactions.
As the analysis progresses to smaller subsets of variables, a clear pattern emerges. The models increasingly focused on core yield components, key quality traits, and essential trial information. This dimensionality reduction suggests that these features are the most critical for accurate potato classification. Overall, this comprehensive examination of variable importance provides valuable insights into the factors that most significantly influence potato classification, offering potential applications in breeding programs and crop management strategies.
As discussed in Section 2.3, we validated our imputation technique by comparing the results of running our predictive models on a complete dataset and on the same dataset with values randomly removed. As with our main results, columns containing 400 or more NA’s were removed, leaving a total of 885 observations, 404 of which were complete. Considering the entire dataset, we computed the distributions of NA’s for each feature. We then randomly removed values from each feature in the 404-observations set according to their NA distributions. Table 4, provides the performance measures obtained from the top three models on each of these sets.
Table 4: Performance measures of the top three models on the complete dataset of 404 observations and the same dataset with randomly removed values following the NA distributions of the entire dataset of 885 total observations.
@@
Model Accuracy F1-score MCC Accuracy F1-score MCC
Neural Net 0.97 0.88 0.766 0.95 0.80 0.623
HGBC 0.94 0.75 0.541 0.93 0.73 0.485
SVM 0.97 0.88 0.766 0.96 0.82 0.692
3.2 Simulation Study Results
Table 5: Performance metrics for original and simulated datasets.
TRIAL MODEL TP FP FN TN ROC MCC SE
original Neural Net 136 10 7 8 0.73242 0.428767
original HGBC 143 3 10 5 0.656393 0.418387
original SVM 142 4 8 7 0.719635 0.506071
original LR 132 14 6 9 0.752055 0.418739
normal_uniform Neural Net 137 9 11 4 0.588172 0.192024 0.00494426
normal_uniform HGBC 139 7 10 5 0.656313 0.346785 0.00438808
normal_uniform SVM 141 5 10 5 0.663721 0.41474 0.0033404
normal_uniform LR 134 12 8 7 0.675231 0.322613 0.00376363
normal_beta Neural Net 137 9 11 4 0.593275 0.202736 0.00505157
normal_beta HGBC 140 6 10 5 0.658075 0.357357 0.00437405
normal_beta SVM 142 4 10 5 0.666871 0.424167 0.00297892
normal_beta LR 133 13 8 7 0.679679 0.328813 0.00386298
gamma_uniform Neural Net 137 9 11 4 0.593 0.202951 0.00510945
gamma_uniform HGBC 139 7 10 5 0.6597 0.357473 0.00440051
gamma_uniform SVM 141 5 10 5 0.664447 0.414572 0.00327735
gamma_uniform LR 133 13 8 7 0.679753 0.328597 0.00384738
gamma_beta Neural Net 137 9 11 4 0.59558 0.209727 0.00525216
gamma_beta HGBC 140 6 10 5 0.654283 0.35311 0.00440901
gamma_beta SVM 142 4 10 5 0.664786 0.417948 0.00310447
gamma_beta LR 133 13 8 7 0.676524 0.32357 0.00396213
The simulation study conducted here offered valuable insights into the performance and robustness of various machine learning algorithms across different data distributions. By comparing the original data performance with results from four simulated distributions, we were able to draw significant conclusions about each model’s strengths and weaknesses. The simulation study results (see Table 5) indicate that non-linear models, such as Support Vector Machine (SVM), Histogram-based Gradient Boosting Classifier (HGBC), and Neural Networks, generally outperform traditional linear models like Logistic Regression in terms of classification metrics such as True Positives (TP), False Positives (FP), True Negatives (TN), False Negatives (FN), area under the receiver operating characteristic curve (ROC score), and Matthews Correlation Coefficient (MCC). These findings highlight the importance of selecting appropriate machine learning models for agricultural trials, where data can be complex and non-linear relationships are prevalent.
Looking at the original data performance, we find that each model has its unique characteristics. The Neural Net achieved high true positives (136) but also had relatively high false positives (10), resulting in the second-highest ROC score (0.73242). The HGBC demonstrated excellent precision with the highest true positives (143) and lowest false positives (3), although it had the lowest ROC score (0.656393) on the original data. The SVM showed a balanced performance with high true positives (142), low false positives (4), and the highest MCC (Matthews Correlation Coefficient) of 0.506071, suggesting good overall performance. Logistic Regression achieved the highest ROC score (0.752055) but at the cost of the highest false positives (14), indicating a potential trade-off between sensitivity and specificity.
When we examine the performance across the simulated distributions (normal_uniform, normal_beta, gamma_uniform, and gamma_beta), several interesting patterns emerge. All models maintained relatively consistent performance across these distributions, which suggests a degree of robustness to different data generation processes. However, we observe that ROC scores for the simulated datasets generally decreased compared to those of the original data, with the Neural Net showing the most significant drop (from 0.73242 to around 0.59). This decrease hints at the possibility that the models may be slightly overfitted to the original data distribution.
The MCC scores also saw a decrease for all models in the simulated data, with the Neural Net experiencing the most substantial reduction (from 0.428767 to around 0.2). Interestingly, the SVM consistently maintained the highest MCC across all simulations (around 0.41-0.42), indicating its strong generalization capabilities. The standard error values remained consistently low across all simulations, ranging from about 0.003 to 0.005. These low values indicate high precision in the performance estimates, lending credibility to the conclusions drawn from this study.
Diving deeper into model-specific performance, the Neural Net showed the largest performance drop in simulated data. While it maintained high true positives (around 137), it saw an increase in false negatives (from 7 to 11), suggesting reduced sensitivity. Its consistently lowest MCC in simulations (around 0.2) further indicates potential overfitting to the original data. This performance drop highlights the importance of rigorous testing across various data distributions when deploying neural networks.
The HGBC, on the other hand, demonstrated consistent performance across simulations. It maintained high true positives (139–140) and low false positives (6–7), suggesting good precision and specificity. The stability of its ROC scores across different distributions (around 0.65–0.66) indicates robustness, making it a reliable choice when the underlying data distribution is uncertain.
The SVM exhibited the most consistent and robust performance across all distributions. It maintained the highest MCC in all scenarios (around 0.41–0.42) and showed a good balance between true positives, false positives, false negatives, and true negatives across simulations. This performance suggests that SVM might be the most reliable choice for this particular problem domain, especially when generalization to unseen data is crucial.
Logistic Regression showed relatively stable performance across simulations and maintained the highest ROC scores in simulated data (around 0.67–0.68), suggesting good discrimination ability. However, it consistently had higher false positives (12–13) compared to other models, indicating a potential bias towards positive predictions. This characteristic could be beneficial or detrimental depending on the specific requirements of the application.
The significance of these results extends beyond the performance of individual models. The study underscores the importance of simulation in assessing model generalization and robustness. The differences observed between original and simulated data performance highlight that evaluating models solely on the original dataset may not provide a complete picture of their real-world performance.
Moreover, the relatively consistent performance across different simulated distributions as demonstrated by the low standard errors, suggests that these models are somewhat invariant to the specific probability distribution of the data. This invariance is a desirable property for many real-world applications where the underlying data distribution may be unknown or subject to change, such as data recorded from potato variety trials conducted in a changing climate.